OpenStudy (rainbow_rocks03):

What is 12/32 in simplest form?

OpenStudy (rainbow_rocks03):

no direct answers plz.

OpenStudy (rainbow_rocks03):

@haleyelizabeth2017

OpenStudy (haleyelizabeth2017):

Start off by dividing both 12 and 32 by 2. What do you get?

OpenStudy (rainbow_rocks03):

12 divided by 2 is 6 and 32 divided by 2 is 16 so my answer is 6/16 right cause that is what I put as my answer

OpenStudy (haleyelizabeth2017):

Nope...they can both be divided by 2 still...

OpenStudy (rainbow_rocks03):

but i didnt submit it yet

OpenStudy (rainbow_rocks03):

ok

OpenStudy (rainbow_rocks03):

so 6/2=3 and 16/2=8 which is 3/8

OpenStudy (haleyelizabeth2017):

Yes!

OpenStudy (rainbow_rocks03):

so that is my answer

OpenStudy (haleyelizabeth2017):

correct.

OpenStudy (rainbow_rocks03):

TY!!! :D

OpenStudy (haleyelizabeth2017):

No problem!

OpenStudy (rainbow_rocks03):

17/4 as a mixed number?

OpenStudy (rainbow_rocks03):

@haleyelizabeth2017

OpenStudy (haleyelizabeth2017):

hmmm...how many times does 4 go into 17?

OpenStudy (haleyelizabeth2017):

Er...what is 16/4?

OpenStudy (rainbow_rocks03):

16/4 is 16 over 4 it is an improper fraction

OpenStudy (rainbow_rocks03):

and 4 can go into 17 4 times

OpenStudy (haleyelizabeth2017):

I know. so, what is 17-16?

OpenStudy (rainbow_rocks03):

is that a minus?

OpenStudy (haleyelizabeth2017):

yes

OpenStudy (rainbow_rocks03):

ok thought so then it is 1

OpenStudy (haleyelizabeth2017):

Right. So we are left with \[4 \frac{1}{4}\]Do you understand how I got that?

OpenStudy (rainbow_rocks03):

not really

OpenStudy (rainbow_rocks03):

kinda though

OpenStudy (haleyelizabeth2017):

Since 4 can go into 17 4 times, that's a whole number. Since it doesn't divide evenly, we have 1 left over (1/4)

OpenStudy (rainbow_rocks03):

the 17-16= the 1 and 4 can go into 17 4 times

OpenStudy (rainbow_rocks03):

ohh

OpenStudy (haleyelizabeth2017):

Yep!

OpenStudy (rainbow_rocks03):

but can you explain the whole thing of how to do it plz?

OpenStudy (rainbow_rocks03):

and TY SOOO much for helping me.

OpenStudy (haleyelizabeth2017):

You're welcome! But that's all there is to explain....16/4=4 17-16 = 1 4 1/4

OpenStudy (rainbow_rocks03):

yeah thats how i wanted to see it...TY again!!!! :D

OpenStudy (haleyelizabeth2017):

:)

OpenStudy (rainbow_rocks03):

just one more plz?

OpenStudy (rainbow_rocks03):

is 6/3 less than more than or equal to 10/15 ?

OpenStudy (rainbow_rocks03):

@haleyelizabeth2017

OpenStudy (rainbow_rocks03):

sorry if i am asking for to much. :(

OpenStudy (haleyelizabeth2017):

multiply 6/3 by 1/5...I think

OpenStudy (haleyelizabeth2017):

or 5/5

OpenStudy (haleyelizabeth2017):

so 30/15? it's larger

OpenStudy (rainbow_rocks03):

so which one is greater then the other?

OpenStudy (rainbow_rocks03):

or are they equal?

OpenStudy (haleyelizabeth2017):

^^^

OpenStudy (haleyelizabeth2017):

look at my last post.

OpenStudy (rainbow_rocks03):

hang on let me write down how the quetion looks

OpenStudy (haleyelizabeth2017):

Okay...

OpenStudy (rainbow_rocks03):



OpenStudy (haleyelizabeth2017):

> I think

OpenStudy (rainbow_rocks03):

you think it is A?

OpenStudy (haleyelizabeth2017):

Yes

OpenStudy (rainbow_rocks03):

cause you only put i think

OpenStudy (rainbow_rocks03):

ok thx for the help

OpenStudy (haleyelizabeth2017):

Yw